Kate wants to prepare a snack mix comprised of peanuts and dried fruit. Peanuts cost $7 a pound and dried fruit costs $5 a pound. Kate wants to spend exactly $45 and prepare 8 pounds of her snack mix. How many pounds of peanuts should Kate purchase, and how many pounds of dried fruit should Kate purchase? Show work.
Answer by mananth(16949)

Peanuts cost $7 a pound . let this be x pounds
and dried fruit costs $5 a pound. this will be 8-x pounds
Mix cost $45 and prepare 8 pounds
...
cost equation
7x+5*(8-x)= 45
2x+40-5x=45
2x=45-40
2x=5
/2
x=5/2 pounds of peanuts
=2.5 pounds
balance 5.5 pounds of fruit
...
m.ananth@hotmail.ca